# FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning

### v2

@@ Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps. To bridge this gap, we propose a question answering dataset for formula-based numerical reasoning called FormulaQA, from junior high school physics examinations.  We further conduct evaluations on LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thoughts methods and we explored the approach of using retrieval-augmented LLMs when providing an external formula database. We also fine-tune on smaller models with size not exceeding 2B. Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaQA.
FormulaQA: A Question Answering Dataset for
Formula-Based Numerical Reasoning
Xiao Li and Sichen Liu and Bolin Zhu and Yin Zhu and Yiwei Liu and Gong Cheng State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,sichenliu,bolinzhu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
@@ 1 Introduction
Dataset Math23K-F MAWPS-F FormulaQA
# questions 23,162 2,373 5,420
# formulas (and variants) 51 (131) 18 (46) 271 (824)
# questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) 5,420 (100%)
Avg. # reasoning steps 1.16 1.01 2.37
Table 1: Statistics of Math23-F, MAWPS-F and our FormulaQA.
Figure 1: An example taken from FormulaQA. All the numerical values (including units) are highlighted in red font. In the formulas, we marked the calculated intermediate results with purple font. All formulas are highlighted in blue.
@@
Numerical reasoning constitutes one of the significant forms within natural language reasoning Frieder et al. (2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs OpenAI (2023); Touvron et al. (2023); Li et al. (2023c) and specialized datasets Wang et al. (2017); Dua et al. (2019); Amini et al. (2019); Cobbe et al. (2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets do not fully address the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination Frieder et al. (2023); Bang et al. (2023). Consequently, one might naturally wonder that “What knowledge could I use to guide numerical reasoning process?”. Formulas just represent the type of knowledge that is often overlooked in research but frequently utilized in real-life applications.
@@
Take a question from the GSM8K Cobbe et al. (2021a) dataset as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in FormulaQA, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object. We annotated formulas for each question in FormulaQA. An example of FormulaQA is shown in Figure 1. The formula-based feature makes FormulaQA a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. In fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider the incorporation of formulas into numerical reasoning.
@@
Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F. However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in each dataset, respectively, require the use of formulas.
@@
To address this gap, we constructed a question answering dataset for numerical reasoning that requires the use of formulas called FormulaQA. We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaQA with two existing formula-based datasets. In comparison to Math23K-F and MAWPS-F, FormulaQA contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaQA require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaQA is more challenging and better suited for evaluating existing models as a multi-step reasoning task.
@@
We evaluated LLMs ranging from 7B to >100B parameters, as well as fine-tuned smaller models such as mT5-base/large Xue et al. (2021) and Qwen-1.8B Bai et al. (2023) on FormulaQA. We also attempted to train an encoder for formula retrieval and experimented with retrieval-augmented generation models. Our evaluation findings show that the best existing models only achieve an accuracy of around 70%, which is not yet satisfactory, indicating that there is still significant room for exploration in formula-based numerical reasoning.
Our contributions can be summarized as follows:
@@
  • We construct a formula-based numerical reasoning dataset FormulaQA, with comprehensive annotations for each question.
@@
  • We conduct evaluations on LLMs of various sizes, fine-tuned smaller models, and retrieval-augmented generative models, establishing a strong baseline for future research.
@@
Dataset and code are available on https://github.com/nju-websoft/FormulaQA under the Apache License 2.0.
@@ 2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP Dua et al. (2019), GSM8K Cobbe et al. (2021b), TSQA Li et al. (2021) and MATH Hendrycks et al. (2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., the question) and require the generation of an arithmetic expression that can compute the answer. Representative datasets include MAWPS Koncel-Kedziorski et al. (2016), Math23K Wang et al. (2017), MathQA Amini et al. (2019), etc.
The recently introduced datasets Liu et al. (2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost=unit_cost $\times$ total_number). Our FormulaQA dataset collects questions from junior high school physics examinations, with each question accompanied by domain-specific formulas. In addition, we have also annotated a formula database for FormulaQA that can serve as an external knowledge base, used to assess retrieval-augmented systems.
@@ 2.2 Numerical Reasoning Methods
The methods for solving numerical reasoning have evolved from statistical approaches Hosseini et al. (2014); Kushman et al. (2014) to those based on rules and templates Shi et al. (2015); Wang et al. (2019) and further to methods based on deep learning models Gupta et al. (2019); Chen et al. (2022); Kim et al. (2022); Li et al. (2023a). In the past two years, with the rapid development of LLMs, LLMs have demonstrated strong capabilities in resolving numerical reasoning questions. Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method Wei et al. (2022), along with many subsequent variant approaches Kojima et al. (2022); Wang et al. (2022); Zhou et al. (2022); Li et al. (2023b).
We have established representative existing methods as baselines for FormulaQA, including fine-tuning on smaller-scale models with fewer than 2 billion parameters, such as mT5-base/large, and Qwen-1.8B. Additionally, we have applied zero-shot and few-shot CoTs prompting methods to LLMs ranging from 7 billion to over 100 billion parameters. Given that existing numerical reasoning methods scarcely consider retrieval-based approaches, as the existing numerical reasoning datasets mostly rely on implicit commonsense knowledge, in FormulaQA, we have trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
@@ 3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present, filtering out those that do not require formulas for solving. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question included an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all parameters within the formulas. Finally, we compiled all the formulas we obtained and merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaQA in detail below.
@@ 3.1 Preprocessing
We crawled 18,433 junior high school physics examination questions in China from 2015 to the present from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions. Each original question contains a question text and an explanation text that includes the reasoning steps. We eliminated questions requiring diagrams.
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a rule-based approach, we extracted the final numeric answer, including its corresponding unit, from the explanation. It was observed that for 487 instances, the final answer had not been accurately extracted from the explanations, prompting us to perform manual computations and annotations to determine the correct final answers for these questions. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
3.2 Formula Normalization
@@
Calculating the degree of temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃. The degree of water temperature increase = 40 ℃.
Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J. The heat absorbed by water = 8400000 J.
The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%. The thermal efficiency of the water heater = 84%.
@@
We found that the reasoning steps (i.e. the explanation) in the obtained raw dataset lacked a normalized format and were expressed quite casually. Some formulas mixed parameter names (e.g., “mass of water”) and symbols (e.g., “$m_{water}$”), while others simply provided calculations in numerical form without parameter names or symbols. In order to ensure that explanations were based on a normalized form of formulas, we then normalized the formula annotations in the explanations. An example can be found in Table 2. In this process, we need to identify the formulas used within the original explanations and to correct any formatting issues. Manually undertaking such tasks required significant effort. However, since the process is not open-ended, but rather structured, methodical and verifiable, we ensured the accuracy and consistency of the numerical reasoning by applying the extraction of formulas from the explanations. Once the formulas were identified, we could calculate each step to ensure that the results aligned with the corresponding answers. Any redundant formulas or missing parameters could be detected, so we could introduce the assistance of LLMs here.
@@
To enhance the efficiency of the annotation process, during the LLM annotation phase, we adopted a coarse-to-fine annotation approach with the help of a LLM [^1] During dataset constructions, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start). Qwen-max is a LLM with over 100 billion parameters and a strong capability in understanding Chinese.. We first prompted the LLM in a few-shot manner to generate accurate explanations of the reasoning process. Then, we used few-shot prompts to guide the LLM in correcting minor errors within the normalized explanations, including formatting errors in formula annotations and inaccuracies in the parameters used during computations. Both prompts can be found in Appendix A.1.1. Next, we will provide a detailed description of this process.
@@
Initially, we introduced the question along with its original explanation and the corresponding answer to guide the LLM through few-shot prompting to revise the original explanation. We observed that the ability of the LLM to revise explanations towards normalized explanations remains satisfactory. To assess the correctness of the revised explanations, we extracted formulas from these explanations and then computed the answer using the numbat tool [^2] https://numbat.dev. Numbat is designed for scientific computations with support for physical dimensions and units.. In addition to providing explanations, we also required the LLM to present the values, symbols, and units of each parameter in the formulas in the form of a table. An example is shown in Figure 1.
@@
At this stage, we used a program to check the correctness of the formula format in the explanations, including whether there were omissions in parameter names, parameter symbols, and corresponding units and these issues are all correctable. Therefore, if our program detected that the LLM had not successfully generated an accurate normalized explanation, we used few-shot prompting to identify and correct these specific errors. More details can be found in Appendix A.1.1. We observed that the questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, unclear question formulation, and so on. Some examples of these questions can be found in Appendix A.1.2. These questions have been removed from our dataset. Following this step, our dataset contains a remaining total of 5,420 questions.

### v3

# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

@@ Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps. To bridge this gap, we construct a dataset for formula-based numerical reasoning called FormulaReasoning, which consists of 5,420 reasoning-based questions.  We employ it to conduct evaluations of LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thought methods, and we further explore using retrieval-augmented LLMs provided with an external formula database associated with our dataset. We also experiment with supervised methods where we divide the reasoning process into formula generation, parameter extraction, and numerical calculation, and perform data augmentation. Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaReasoning.
@@ 1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Li et al., 2023c) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets have not fully addressed the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination (Frieder et al., 2023; Bang et al., 2023). Consequently, one might naturally ask “What knowledge could I use to guide numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in research but is frequently utilized in real-life applications.
Figure 1: An example taken from FormulaReasoning. Numerical values (including units) given in the question and obtained from intermediate steps are highlighted in red and purple, respectively. Formulas and their elements are in blue.
@@
Take a question from the GSM8K (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in our FormulaReasoning dataset, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object.
@@
Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F. However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in these datasets, respectively, require the use of formulas.
To address this gap, we constructed a dataset for numerical reasoning that requires the use of formulas called FormulaReasoning. We annotated formulas for each question in FormulaReasoning. An example of FormulaReasoning is shown in Figure 1.111Please note that FormulaReasoning is in Chinese. For the convenience of understanding, we translated Chinese into English in all the examples presented in this paper. The formula-based feature makes FormulaReasoning a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. Indeed, in fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider explicit incorporation of formulas into numerical reasoning.
Dataset Math23K-F MAWPS-F GSM8K FormulaReasoning # questions 23,162 2,373 8,792 5,420 # formulas (and variants) 51 (131) 18 (46) 0 (0) 272 (824) # questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) N/A 5,420 (100%) Avg. # reasoning steps 1.16 1.01 3.59 2.37
Table 1: Statistics of Math23-F, MAWPS-F, GSM8K and our FormulaReasoning.
@@
We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and a set of relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known GSM8K. In comparison to Math23K-F and MAWPS-F, FormulaReasoning contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaReasoning require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaReasoning is more challenging and better suited for evaluating existing models as a multi-step formula-based reasoning task.
@@
We used FormulaReasoning to evaluate LLMs ranging from 7B to >100B parameters, as well as fine-tuned models such as Qwen-1.8B (Bai et al., 2023) and ChatGLM-6B (Zeng et al., 2022) with a proposed Chain-of-Thought supervised fine-tuned method and a data augmentation method. We also trained an encoder for formula retrieval and experimented with retrieval-augmented generative models. Our empirical findings show that the best existing models only achieve an accuracy of around 74%, lagging behind an accuracy 92% of humans, indicating that there is still significant room for exploration in formula-based numerical reasoning.
Our contributions are summarized as follows:
@@
  • We construct a formula-based numerical reasoning dataset FormulaReasoning, with fine-grained annotations for each question. As a formular knowledge-guided numerical reasoning dataset, it can be applied to tasks involving trustworthy and verifiable reasoning.
@@
  • We conduct evaluations on LLMs of various sizes, supervised fine-tuned models, and retrieval-augmented generative models. The experimental results establish a strong baseline for future research and also indicate that the task remains unresolved.
@@
The dataset is available on https://zenodo.org/doi/10.5281/zenodo.11408109 under the CC BY 4.0 License and our code is available on https://github.com/nju-websoft/FormulaReasoning under the Apache License 2.0.
@@ 2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP (Dua et al., 2019), GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021) and MATH (Hendrycks et al., 2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc.
The recently introduced datasets (Liu et al., 2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number). By contrast, our FormulaReasoning dataset collects questions from junior high school physics examinations, with every question accompanied by formulas. In addition, we also annotated a formula database for FormulaReasoning that can serve as an external knowledge base, used to assess retrieval-augmented systems.
@@ 2.2 Numerical Reasoning Methods
The methods for solving numerical reasoning have evolved from statistical approaches (Hosseini et al., 2014; Kushman et al., 2014) to those based on rules and templates (Shi et al., 2015; Wang et al., 2019) and further to methods based on deep learning models (Gupta et al., 2019; Chen et al., 2022; Kim et al., 2022; Li et al., 2023a). In the past two years, with the rapid development of LLMs, LLMs have demonstrated strong capabilities in resolving numerical reasoning questions. Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method (Wei et al., 2022), along with many subsequent variant approaches (Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022; Li et al., 2023b).
We established representative existing methods as baselines for FormulaReasoning, including zero/few-shot CoTs prompting methods to LLMs ranging from 7B to over 100B parameters. We trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning. We also divided the reasoning process into formula generation, parameter extraction, and calculation, and used data augmentation to enhance fine-tuned models with fewer than 7B parameters.
@@ 3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present. We had a total of five postgraduate volunteer students, and they all hold a bachelor’s degree in science and engineering. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question is associated with an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all the parameters within the formulas. Finally, we compiled all the formulas we merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaReasoning in detail below.
@@ 3.1 Preprocessing
We crawled 18,433 junior high school physics examination questions in China from 2015 to the present from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions. Each raw question contains a question text and an explanation text that includes the reasoning steps. We eliminated questions requiring diagrams.
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a regular expression-based approach, we extracted the final numeric answer, including its unit, from the explanation. We found that for 487 questions, the regular expressions did not return results, so we manually annotated the positions of their answers in the text explanations. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
@@
1. Calculating the temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃. The degree of water temperature increase = 40 ℃.
2. Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J. The heat absorbed by water = 8400000 J.
3. The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%. The thermal efficiency of the water heater = 84%.
@@ 3.2 Formula Normalization
We found that the reasoning steps (i.e. the explanation) in the obtained raw dataset lacked a normalized format and were expressed quite casually. Some formulas mixed parameter names (e.g., “mass of water”) and symbols (e.g., “$m_{water}$”), while others simply provided calculations in numerical form without parameter names or symbols. In order to ensure that all explanations adopted a normalized form of formulas, we normalized the formula annotations in the explanations. An example can be found in Table 2. In this process, we need to identify the formulas used within the original explanations and to correct any formatting issues. Manually undertaking such tasks would require significant effort. However, since the process is not open-ended, but rather structured and verifiable, we could automatically, e.g., using a LLM, extract formulas from the explanations, calculate each step, and compare the result with the given answer to ensure the accuracy of this normalization process.
@@
Specifically, to enhance the efficiency of the annotation, we adopted a coarse-to-fine annotation approach with the help of a LLM [^2] During dataset construction, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start). Qwen-max is a LLM with over 100B parameters and a strong capability in Chinese.. We first prompted the LLM in a few-shot manner to generate accurate explanations of the reasoning process. Then, we used few-shot prompts to guide the LLM in correcting minor errors within the normalized explanations, including formatting errors in formula annotations and inaccuracies in the parameters used during computations. Both prompts can be found in Appendix A.1.1. Next, we will provide a detailed description of this process.
@@
Initially, we introduced the question along with its original explanation and the corresponding answer to guide the LLM through few-shot prompting to revise the original explanation. We observed that the ability of the LLM to revise explanations towards normalized explanations remained satisfactory. To assess the correctness of the revised explanations, we extracted formulas from these explanations and then computed the answer using the numbat tool [^3] https://numbat.dev. Numbat is designed for scientific computations with support for physical units.. In addition to providing explanations, we also required the LLM to present the values, symbols, and units of each parameter in the formulas in the form of a table. An example is shown in Figure 1.
@@
At this stage, we checked the correctness of the formula format in the explanations by automatic rules, including whether there were omissions in parameter names, parameter symbols, or corresponding units, and these issues were all correctable. Therefore, if our program detected that the LLM had not successfully generated an accurate normalized explanation, we used few-shot prompting to identify and correct these specific errors. More details can be found in Appendix A.1.1. We observed that the questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, unclear question formulation, and so on. Some examples of these questions can be found in Appendix A.1.2. These questions were removed from our dataset. Following this step, our dataset contains a remaining total of 5,420 questions.